OpenStudy (anonymous):

how does cos2x come out to 1-2sin^2x?

OpenStudy (inkyvoyd):

Well...

OpenStudy (inkyvoyd):

cos(a+b)=cos a cos b-sin a sin b

OpenStudy (inkyvoyd):

cos(theta+theta)= cos theta cos theta - sin theta sin theta

OpenStudy (inkyvoyd):

thus cos(2theta)=cos^2 theta - sin^2 theta

OpenStudy (inkyvoyd):

we know that sin^2 theta+ cos^2 theta=1 for all real values of theta

OpenStudy (inkyvoyd):

so we can safely substitute cos^2 theta=1- sin^2 theta

OpenStudy (anonymous):

wowww thank you soo much! haha that makes a lot of sense! :)

OpenStudy (inkyvoyd):

kay :) Just a tip here, whenever you see sin^2 or cos^2 x always see if you can simplify it using substitution. Never forget that sin^2 x+cos^2 x=1!

OpenStudy (anonymous):

ok!